How many Liters of a 60% acid solution must be mixed with a 75% acic solution to get 20 Liters of a 72% solution?
Here is what I have tried. But it does not make sense.
x=60% y=75%
x+y=20L
60%x+75%y=72%
I then tried to solve for x and y and the answers were x=19.04 y=.96 I do not know what units.
I use to be good at stuff like this, but I have not done it since a year ago. Please help me.

Answer by fractalier(2101)

Let x be the amount of 60% acid. The setup is like this:
60x + 75(20-x) = 72(20)
60x + 1500 - 75x = 1440
-15x = -60
x = 4
Thus 4 L of 60% and 16 L of 75% acid...